OpenStudy (samigupta8):

The plane containing the line x-1/1=y-2/2=z-3/3 and parallel to the line x/1=y/1=z/4 passes through the point A.(1,0,5) B.(0,3,-5) C.(-1,-3, 0) (1,-2, 5)

OpenStudy (michele_laino):

here is my reasoning: the general equation of a plane is: \[ax + by + cz = d,\quad a,b,c,d \in \mathbb{R}\]

OpenStudy (michele_laino):

so, we have to compute the parameters \(a,\;b,\;c,\;d\)

OpenStudy (michele_laino):

from general theory, we know that the vector \((a,b,c)\), is orthogonal with respect to the plane: 

OpenStudy (samigupta8):

Yep....

OpenStudy (michele_laino):

next: the parametric equations of parallel line, are: \[\left\{ {\begin{array}{*{20}{c}} {x = t} \\ {y = t} \\ {z = 4t} \end{array}} \right.\] so, its direction vector is: \((1,1,4)\) and we can write: \[\left( {a,b,c} \right) \cdot \left( {1,1,4} \right) = 0 \Rightarrow a + b + 4c = 0\;\left( * \right)\] 

OpenStudy (samigupta8):

Yep right...

OpenStudy (michele_laino):

next: the contained line has the subsequent parametric equations: \[\left\{ {\begin{array}{*{20}{c}} {x = 1 + t} \\ {y = 2t + 2} \\ {z = 3t + 3} \end{array}} \right.\]

OpenStudy (michele_laino):

the contained line does belong to the plane, if and only if, the subsequent conditions hold: \[a\left( {1 + t} \right) + b\left( {2t + 2} \right) + c\left( {3t + 3} \right) = d\]

OpenStudy (michele_laino):

from which I get: \[\left( {a + 2b + 3c} \right)t + a + 2b + 3c = d\]

OpenStudy (michele_laino):

therefore, I can write: \[\left\{ \begin{gathered} a + 2b + 3c = 0 \hfill \\ \hfill \\ a + 2b + 3c = d \hfill \\ \end{gathered} \right.\] which tell us that \(d=0\), and: \[a + 2b + 3c = 0\;\left( {**} \right)\]

OpenStudy (michele_laino):

now, I have to solve the system of equations \((*)\) and \((**)\): \[\left\{ \begin{gathered} a + b + 4c = 0 \hfill \\ a + 2b + 3c = 0\; \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

one of the corresponding solution (please note that we have \(\infty\) solutions), is: \[a = - 5,\;b = 1,\;c = 1\]

OpenStudy (michele_laino):

and the requested plane is: \[ - 5x + y + z = 0\]

OpenStudy (michele_laino):

so, what can you conclude?

OpenStudy (samigupta8):

First option....:)

OpenStudy (michele_laino):

that's right!! :)

OpenStudy (samigupta8):

Thanks sir!!!

OpenStudy (michele_laino):

:)